### v3

# Let’s Fuse Step by Step: A Generative Fusion Decoding Algorithm with LLMs for Multi-modal Text Recognition

Abstract
We introduce “Generative Fusion Decoding” (GFD), a novel shallow fusion framework, utilized to integrate Large Language Models (LLMs) into multi-modal text recognition systems such as automatic speech recognition (ASR) and optical character recognition (OCR). We derive the formulas necessary to enable GFD to operate across mismatched token spaces of different models by mapping text token space to byte token space, enabling seamless fusion during the decoding process. The framework is plug-and-play, compatible with various auto-regressive models, and does not require re-training for feature alignment, thus overcoming limitations of previous fusion techniques. We highlight three main advantages of GFD: First, by simplifying the complexity of aligning different model sample spaces, GFD allows LLMs to correct errors in tandem with the recognition model, reducing computation latencies. Second, the in-context learning ability of LLMs is fully capitalized by GFD, increasing robustness in long-form speech recognition and instruction aware speech recognition. Third, GFD enables fusing recognition models deficient in Chinese text recognition with LLMs extensively trained on Chinese. Our evaluation demonstrates that GFD significantly improves performance in ASR and OCR tasks, with ASR reaching state-of-the-art in the NTUML2021 benchmark. GFD provides a significant step forward in model integration, offering a unified solution that could be widely applicable to leveraging existing pre-trained models through step by step fusion.  [^1] Code is available at https://github.com/mtkresearch/generative-fusion-decoding
1 Introduction
In recent years, integrating Large Language Models (LLMs) into multi-modal systems has emerged as a frontier, offering substantial advancements in applications such as automatic speech recognition (ASR) [40], visual question answering (VQA) [36], reinforcement learning [50], and generalist agent [42]. The robust language model capabilities that the LLM offers make it an ideal candidate for integration with other models, many of which are “undertrained” in the textual domain, compared to trillions of tokens consumed by the LLMs. Obtaining sufficient quality paired data and training resources in a huge endeavor for text recognition systems such as automatic speech recognition (ASR) and optical character recognition (OCR).
The integration of heterogeneous models has been explored extensively in ASR literature, as exemplified by strategies such as shallow fusion, late fusion, mid fusion, and early fusion. Works in shallow fusion [9, 27, 10] interpolates text recognition outputs with LM outputs at each step of beam search. Late fusion approaches typically engage LLMs to score the initial sentence generated by the recognition model. The scores are then utilized for fusing with recognition probabilities for joint transcription [6], for re-ranking the the constrained set of hypothesis as work in re-scoring [46], for generating new hypothesis options through Generative Error Correction (GER) [5]. Notably, methods in mid-fusion propose fusing models through model attentions to integrate LLM into the decoding process of the recognition model [41, 37], and work in early-fusion has proposed concatenating speech embeddings or image embeddings along with LLM embedings as inputs to the LLM for ASR and VQA [17, 7]. However, the aforementioned fusion methods bear several challenges. The early fusion techniques are often limited by a low correction rate due to the restricted range of available choices for re-evaluation. Mid-fusion discards the whisper decoder, which has been shown to have denoising ability[19]. Additionally, late fusion with GER may lead to hallucination due to the LLM as the LLM generation is devoid of auditory inputs. Due to the requirements for shallow and late fusion to align output spaces, and the need for mid and early fusions to involve joint training with additional paired data, prevent an LLM from being used as is with existing text-recognition models without proper feature alignment through trainings.
Figure 1: The GFD integrated framework. The framework aims to integrate pre-trained text-recognition models with LLMs to augment the recognition capabilities. However, direct fusion is impeded by the discrepancy in token spaces between the two model types. To overcome this obstacle, GFD is introduced, which converts the token space into a byte space, thereby facilitating incremental fusion. Equations 7 and 8 are employed to compute the fusion probability, enabling a stepwise fusion process. The symbols $\mathcal{Z}^{\texttt{TR}}$ and $\mathcal{Z}^{\texttt{LM}}$ denote the contextual information pertinent to the text-recognition models and the language models, respectively. Notably, $\mathcal{Z}^{\texttt{TR}}$ may correspond to either an audio input for ASR or a text image for OCR, whereas $\mathcal{Z}^{\texttt{LM}}$ typically represents a textual prompt.
To address the aforementioned challenges, we introduce a novel shallow fusion framework, termed “Generative Fusion Decoding” (GFD). GFD resolves the token space mismatch between models by bridging heterogeneous spaces through the expected score values of plausible sequences and mapping text token space to byte token space. This allows GFD to seamlessly fuse LLMs with text recognition models during the decoding process. Our unified framework is plug-and-play, applicable to text recognition tasks such as ASR and OCR, and compatible with diverse auto-regressive text recognition models as well as language models. We highlight three main advantages of GFD. First, GFD addresses key limitation of previous fusion techniques — namely the complexity of aligning different model sample spaces via re-training associated with prior deep integration [41] or two-pass [46] strategies. It allows LLMs to participate in the sampling process as conventional shallow fusion, providing real-time correction of errors generated by the text recognition model, thereby extending the influence of LLMs beyond the complete outputs of the recognition model. Secondly, as modern LLMs are trained on trillions of multilingual text tokens [22, 26], generally around 3 orders of magnitude more than end-to-end ASR or OCR models [40], LLMs in GFD can be seen as a strong scoring model in the step by step decoding that improves the multi-lingual outputs of ASR and OCR systems. Lastly, because of the in-context learning ability of LLMs, the GFD fusion system with LLMs and ASR can readily leverage domain information as prompts effectively for domain adaptation similarly to prior work through re-scoring [14] or instruction tuning [35].
To evaluate the efficacy of our algorithm, we implemented the GFD for ASR and OCR tasks. In the ASR task, we evaluated on the conventional ASR task as shown in Section 4.2 for English and Chinese audios. The evaluation of Chinese ASR is motivated by existing literature [40], which indicates a performance outlier in Chinese compared to other languages despite abundant paired-training data. We hypothesized that LLMs could enhance Chinese ASR by offering additional scoring to differentiate between semantically distinct but homophonous words. Furthermore, we demonstrate that ASR with GFD can improve recognition for long-form audio through leveraging the long-context ability of LLMs (Section 4.3); for context-aware recognition for rare words and domain tags via in-context learning capabilities of LLMs (Section 4.4). Lastly, application to the OCR task, particularly targeting long text OCR for historical documents, demonstrates that OCR with GFD can significantly improve recognition accuracy (Section 4.5).
2 Related work
2.1 Model fusion
Training a multi-objective model from scratch is often costly [3, 2, 1, 15]. Consequently, researchers have pivoted towards combining existing models with different modalities to improve accuracy without the prohibitive costs of building new systems from the ground up.
Model fusion developed in the field of ASR provides a plausible path for combining existing trained models. The technique have evolved significantly in recent years, encompassing a variety of approaches designed to integrate different models to enhance performance. Shallow fusion or late fusion, often employed in ASR, combines end-to-end ASR models with external language models at the decoding level, improving recognition accuracy without altering the underlying ASR architecture [27, 25, 6, 52]. However, due to the heterogeneous sample spaces of models, the prerequisite of shallow and late fusion requires aligning sample spaces of model distributions, enabled through fine-tuning a projection module [6]. Late fusion training methods may suffer from modality laziness problem in tasks where uni-modal priors are meaningful [16]. Work in re-scoring N-best hypothesis from the first pass inference with an external language model has been shown effective in reducing recognition errors [43, 23, 46]. However, the method prunes the correct prefix early on away from N-best list and requires a second pass in the decoding process.
Alternatively, deep fusion integrates models at the level of hidden feature, requiring fine-tuning models to fuse deep features [20]. Cross-modal fusion, similar to deep fusion, integrates pre-trained end-to-end ASR model with LLM [41, 51, 33] or vision model with LLM [7, 36] via learning a joint representation with large amount of extra paired audio-text or image-text data.
Our newly proposed approach—GFD—decodes in homogeneous sequence element space with byte strings, and therefore relaxes the alignment of sample spaces of distributions without any model training. Without pre-training or fine-tuning, we ensure that the individual pre-trained model performance is guaranteed on their respective task, and is not influenced by volatility of further training.
2.2 Contextual conditioning
Auto-regressive LLMs have exhibited capabilities in in-context learning [39], instruction following [38], and knowledge synthesis [44]. Such capabilities have been applied to solving domain adaptation in speech recognition for rare words or out-of-domain context through contextual biasing [10] and prompting fine-tuned models [35, 47, 33, 47]. With GFD, such problem setting can be tackled by leveraging a performant LLM as is through prompting without further fine-tuning.
2.3 Chinese ASR
One of the most significant challenges in developing ASR systems for the Chinese language stems from its highly homophonous nature [30, 29, 8]. Unlike English, where there is a larger variety of phonemes and a relatively consistent correspondence between spelling and sound, Chinese relies on a limited set of tones and syllables to represent thousands of characters. Consequently, Chinese ASR systems must not only accurately capture the tonal nuances but also analyze the linguistic context to disambiguate these homophones. This requires sophisticated LMs that can effectively leverage contextual clues to predict the intended characters. The integration of LLMs has shown promise in addressing these challenges [11, 31, 34]. GFD is also expected to enhance the performance of Chinese ASR by utilizing the contextual conditioning capabilities of LLMs.
3 Method
3.1 Generative fusion decoding
For conditional text generation models, the sequence with the highest probability during inference is found using the following formula:
$\{T_{s}\}^{*}=\underset{\{T_{s}\}}{\mathrm{arg\,max}}\log\mathbb{P}(\{T_{s}\},\mathcal{Z}),$ (1)
where $\{T_{s}\}$ represents the sequence of tokens generated by the model, and $\mathcal{Z}$ represents the given context or conditioning information, such as audio for speech recognition models [40], images for vision-language-models [1], and prompts for typical language models [4].
Auto-regressive generation is one approach to realize conditional text generation. In this approach, the auto-regressive model is conditioned on the previously generated tokens to generate the next token sequentially. Therefore, the probability $\log\mathbb{P}(\{T_{s}\}|\mathcal{Z})$ is typically decomposed using the chain rule of probability, as follows:
$\log\mathbb{P}(\{T_{s}\},\mathcal{Z})=\sum_{s=1}^{S}\log\mathbb{P}(T_{s}\mid T_{<s},\mathcal{Z}),$ (2)
where $T_{s}$ is the token at position $s$ in the sequence, and $T_{<s}$ represents all the tokens preceding position $s$. In real-world applications, it is impracticable to enumerate all possible token sequences, so beam search is typically used as an approximation method.
In the setting of shallow fusion, we attempt to let multiple models jointly determine the sequence, as shown in the following formula rewrote from Equation (1):
$\{T_{s}^{\texttt{fuse}}\}^{*}$ $=\underset{\{T_{s}^{\texttt{fuse}}\}}{\mathrm{arg\,max}}\sum_{m}\lambda_{m}\log\mathbb{P}_{m}(\{T_{s}^{(m)}\}=\{T_{s}^{\texttt{fuse}}\},\mathcal{Z}^{(m)})$ (3)
where $\{T_{s}^{\texttt{fuse}}\}$ represents the fused sequence of tokens generated by combining the outputs of multiple models, $\lambda_{m}$ is a weighting factor for the $m$-th model, $p_{m}$ denotes the probability distribution of the $m$-th model and $\mathcal{Z}^{(m)}$ represents the context or conditioning information specific to the $m$-th model.
When the sample spaces of models are the same, Equations (3) and (2) can be combined to realize incremental fusion. If the models have different sample spaces due to a mismatch in token spaces, there is no simple way to achieve incremental fusion. One alternative method to approximate Equation (3) is to fuse at the level of the fully generated results from each model. Nevertheless, in practice, fusion at the level of fully generated results poses the problem of an enormous search space because different conditioning variables $\mathcal{Z}^{(m)}$ may produce vastly different results.
To address these challenges, we have introduced a probability transformation, denoted as $\mathcal{M}^{(m)}$, that converts token-level representations into byte-level representations:
$\mathcal{M}^{(m)}:\mathbb{P}_{m}(\{T_{s}\},\mathcal{Z}^{(m)})\longrightarrow\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)}),$ (4)
where $\{B_{l}\}$ represents the sequence of bytes after the transformation, and $l$ denotes the position in the byte sequence. This transformation allows for a unified representation across different models, facilitating the fusion process even when the original token spaces differ. The byte-level fusion can then be performed using a similar approach to Equation (3), but with the byte-level probabilities:
$\{B_{l}^{\texttt{fuse}}\}^{*}$ $=\underset{\{B_{l}^{\texttt{fuse}}\}}{\mathrm{arg\,max}}\sum_{m}\lambda_{m}\log\mathbb{P}_{m}(\{B_{l}\}=\{B_{l}^{\texttt{fuse}}\},\mathcal{Z}^{(m)}).$ (5)
To operationalize the probability transformation $\mathcal{M}^{(m)}$, we define a mapping from the token-level probabilities to the byte-level probabilities. This mapping takes into account the prefix relationship between the token sequence and the byte sequence. Specifically, we express the byte-level probability $\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})$ as a sum over all possible token sequences that share a common prefix with the byte sequence $\{B_{l}\}$. The probability of each token sequence is computed as the product of the conditional probabilities of each token given the preceding tokens and the context $\mathcal{Z}^{(m)}$. This relationship is formalized in the following equation:
$\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})=\sum_{\{B_{l}\}}\biggr{[}\prod_{s}\mathbb{P}_{m}(T_{s}^{(m)}\mid T_{<s}^{(m)},\mathcal{Z}^{(m)})\biggr{]}_{\{T_{s}\}}\times\mathbb{1}\bigl{(}\{T_{s}\}\texttt{.prefix}=\{B_{l}\}\bigl{)},$ (6)
where .prefix is a function that maps a token sequence to its corresponding byte sequence prefix, and $\mathbb{1}$ is the indicator function that is 1 if the token sequence has the specified byte sequence as a prefix and 0 otherwise. The indicator function ensures that only those token sequences that can be mapped to the given byte sequence contribute to the byte-level probability. In Equation (6), the complexity remains high at $\mathcal{O}(V^{S})$, where $V$ represents the vocabulary size of tokens and $S$ is the sequence length, for identifying token sequences that match the specified criteria of the indicator.
Figure 2: Example of the main sequence and alternative tokens. Assume that the byte sequence is "abcdefghijk". The main token sequence is the tokenization result of the byte sequence and is denoted as $\{\hat{T}_{s}\}$. The alternative tokens are denoted as $T^{\prime}_{i}$.
To reduce complexity, we posit that the main token sequence and its branching alternatives, as shown in Figure 2, dominate the probability contribution, where the main token sequence is produced by applying model tokenization on the byte string and the alternative tokens are essentially look-aheads for potential main tokens that may emerge as the decoding progresses. The significance of the main token sequence is justified by its alignment with the model’s inputs during the original pretraining phase; any other slicing method is penalized in terms of probability due to its lack of representation in the training data. Based on this assumption, we can narrow down the search for token sequences that meet the criteria of $\{T_{s}\}\texttt{.prefix}=\{B_{l}\}$ to only the main token sequence and its branching alternative tokens. We define the main token sequence as $\{\hat{T}_{s}\}$. Given this simplification, we can approximate the byte-level probability $\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})$ by considering only the main token sequence and its immediate alternatives that share the same prefix with the byte sequence $\{B_{l}\}$. This approximation significantly reduces the computational complexity to $\mathcal{O}(V\times S)$ and is expressed in the following equation:
$\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})\approx$  $\texttt{~{}}\mathbb{P}_{m}(\hat{T}_{1}\mid\hat{T}_{<1},\mathcal{Z}^{(m)})\times\bigl{[}$ (7)
 $\texttt{~{}~{}~{}}\mathbb{P}_{m}(\hat{T}_{2}\mid\hat{T}_{<2},\mathcal{Z}^{(m)})\times\bigl{[}$
[…]
 $\texttt{~{}~{}~{}~{}~{}~{}~{}}\mathbb{P}_{m}(\hat{T}_{S}\mid\hat{T}_{<S},\mathcal{Z}^{(m)})$
 $\texttt{~{}~{}~{}~{}~{}~{}~{}}+\sum_{t}\mathbb{P}_{m}(t\mid\hat{T}_{<S},\mathcal{Z}^{(m)})\cdot\mathbb{1}\bigl{(}\{\hat{T}_{<S},t\}\texttt{.prefix}=\{B_{l}\}\bigl{)}$
[…]
 $\texttt{~{}~{}~{}}\bigl{]}+\sum_{t}\mathbb{P}_{m}(t\mid\hat{T}_{<2},\mathcal{Z}^{(m)})\cdot\mathbb{1}\bigl{(}\{\hat{T}_{<2},t\}\texttt{.prefix}=\{B_{l}\}\bigl{)}$
 $\texttt{~{}}\bigl{]}+\sum_{t}\mathbb{P}_{m}(t\mid\hat{T}_{<1},\mathcal{Z}^{(m)})\cdot\mathbb{1}\bigl{(}\{\hat{T}_{<1},t\}\texttt{.prefix}=\{B_{l}\}\bigl{)},$
where $t$ represents an alternative token that could follow the sequence $\hat{T}_{<s}$ and meet the criteria of $\{T_{s}\}\texttt{.prefix}=\{B_{l}\}$. Eventually, substituting Equation (7) into (5) can realize generative fusion decoding (GFD).
In summary, our proposed method for conditional text generation, GFD, through late fusion and byte-level probability transformation offers a novel way to integrate the outputs of multiple models with different token spaces. By transforming token-level probabilities to byte-level probabilities and focusing on the most probable token sequences, we can efficiently fuse model outputs.
3.2 Fusing text-recognition models with LLM
To evaluate the efficacy of our algorithm, we implemented GFD for ASR and OCR tasks. This is achieved by fusing pre-trained text-recognition models with LLMs to enhance recognition capability. In such scenarios, the proposed tokens of ASR and OCR model are conditioned on a modal signal $\mathcal{Z}^{\texttt{TR}}$ while that of the LLM is not. Therefore, the same next-byte probability estimation from the LLM and text recognition model might stem from different token choices. Although the difference would converge eventually as the sequence rolls out, given limited beams in the beam search process, the potential overestimation of implausible words due to scores contributed by LLM spikes the probabilities of common word-leading bytes such as ’t’ and ’r’, and produces locally unstable rankings that may result in false eliminations of correct beams in decoding steps. In response to this challenge, we have adopted text recognition models (ASR and OCR) to propose the next byte/token, while employing LLMs to provide delayed corrective feedback on past tokens. The fusion decoding methodology we have adopted, derived from Equation (5), is delineated in the following formula:
$\{B_{1},..,B_{t}\}^{*}=\underset{\{B_{1},..,B_{t}\}}{\mathrm{arg\,max}}\bigl{[}(1-r)\cdot\log\mathbb{P}^{\texttt{TR}}(\{B_{1},..,B_{t}\},\mathcal{Z}^{\texttt{TR}})$ (8)
$+r\cdot\log\mathbb{P}^{\texttt{LM}}(\{B_{1},..,B_{t-k}\},\mathcal{Z}^{\texttt{LM}})\bigl{]},$
where $\{B_{1},\ldots,B_{t}\}^{*}$ represents the optimal sequence of bytes up to and including position $t$, $p^{\texttt{TR}}$ and $p^{\texttt{LM}}$ are the probability distributions of the text-recognition models and the language models, respectively, $\mathcal{Z}^{\texttt{TR}}$ and $\mathcal{Z}^{\texttt{LM}}$ are the contextual information for the text-recognition and language models, respectively, $r$ is a weighting factor that balances the influence of the text-recognition models and the language models, and $k$ depends on the length of the byte/token proposal. As shown in Figure 1, $\mathcal{Z}^{\texttt{TR}}$ can be either an audio or a text image, and $\mathcal{Z}^{\texttt{LM}}$ can be a prompt.
4 Experiment
4.1 Experimental setup
We evaluate the application of GFD to ASR and OCR tasks. For ASR task, we benchmark datasets in English, Chinese, and Cantonese. The deliberate choice of Chinese and Cantonese is due to that the languages are homophonous and tonal, challenging for ASR systems. This complexity is corroborated by the Fleurs experiment in Whisper [40], where the Chinese word error rate is well above the regressed word error rate in comparison to evaluations in other language at the same amount of pre-training data. Of all tested languages in Whisper, it is the only large-scale language (> 10k hours audio) with such a phenomenon. For OCR, we benchmark image dataset containing long sequence of text as we hypothesize that LLM provide semantic information to an OCR model of recognizing long text sequences.
In terms of the evaluated models, we choose Whisper-large-v2 as the ASR model for greedy and beam search. For our proposed GFD, we use Mistral [26] for English datasets and Breeze [22] for Chinese and Cantonese datasets. We term the GFD fused model with Whisper and Mistral as GFD-ASR-EN; that with Whisper and Mistral as GFD-ASR-ZH. In addition, we benchmark GER, based on task-activating prompting method [5, 48], with Mistral-Instruct and Breeze-Instruct models without finetuning. We include two oracle word error rates following previous work[24], where the N-Best Oracle $o_{nb}$ denotes the error rate calculated with the best candidate in the N-best list, and the Compositional Oracle $o_{cp}$ is the best achievable word error rate using all tokens N-best the list. As to OCR, we utilize TrOCR [32], a pre-trained OCR model, and denote the fused model GFD-OCR.
We benchmark models on various datasets, including Librispeech [21], Medical [18], ATCO2 [45], Fleurs [12], NTUML2021[49], and FormosaSpeech  [^2] https://huggingface.co/datasets/MediaTek-Research/formosaspeech for ASR; NAF [13] for OCR. Librispeech is a collection of corpus from audiobooks with subsets Clean and Other. The Librispeech-Noise is its noised variant, featuring various signal-to-noise ratios, ideally for testing ASR system’s robustness to noise. Medical dataset contains 8.5 hours of medical conversations with associated symptom tags to each audio-text pairs. ATCO2 contains audios of air traffic control communication and accompanied meta-information of airports. Fleurs is a multilingual speech corpus. We deliberately choose Cantonese subset for evaluation as the language is homophonous and tonal. NTUML2021 corpus consists of lecture recordings from the “Machine Learning” course at National Taiwan University in 2021, with corresponding transcriptions and English translations labeled by over 20 bilingual native Chinese speakers. FormosaSpeech corpus includes Chinese recordings of Taiwanese accents amassing up to 6.4 hours of audio-text pairs. NAF consists of images from U.S national archives with labelled bounding boxes and annotations, ideal for evaluating OCR performance.
For all ASR experiments, we report word error rate (WER) for evaluations on English datasets, and mix error rate (MER) for that on Chinese datasets. For OCR experiments, we report character error rate (CER) and exact match (EM). For exact match, we require that the recognized texts fully matches the ground truth. All GFD fused models can be run on a single A6000 GPU. We fix $r=0.2$ in Equation 8 for all our experiments; number of beams equaling 5 and 10 for ASR and OCR experiments, respectively.
4.2 Conventional speech recognition
We first verify the efficacy of GFD in conventional speech recognition setting and report results in Table 1. We notice that beam search improves consistently upon greedy search, and use beam search as the baseline. We found that GFD moderately improves on the Medical and Fleurs-HK dataset. In the more challenging NTUML2021, we obtained a state-of-the-art 8.83 mixed error rate, surpassing prior work[49] and even the oracle N-Best score. This improvement is mostly preserved even when we narrow the target space to only consider English terms in the dataset (Table 2), making GFD-ASR an elegant solution for code-switched ASR. Examples in the Appendix show that improvement can be attributed to the LLM correcting grammatical mistakes and domain terminologies. Performance across Librispeech reveals that GFD improves most on moderate noise conditions but diminishes when the noise is too large $(S/R=-5)$.
In addition, we do not find improvements in the GER setting using non-GER-specific instruct models, consistent with previous work[5]. As we already patched malformatted outputs with the best beam result, this increase of error rate is mostly attributed to LLM hallucinations, including incorrect deletions. We therefore posit that GFD is more robust for incorporating off-the-shelf LLMs on the ASR task than GER.
Table 1: Results on conventional speech recognition.
Whisper GER GFD-ASR-EN Oracle
EN dataset Greedy 5-beams (Baseline) $\text{O}_{\text{nb}}$ $\text{O}_{\text{cp}}$
Librispeech-Clean 2.30 2.28 3.21 2.29 1.91 1.63
Librispeech-Other 5.23 4.97 6.57 4.99 4.20 3.43
Medical 7.30 7.22 7.49 6.74 6.17 5.05
Librispeech-Noise ($S/R=10$) 3.50 3.12 4.05 3.02 2.32 1.94
Librispeech-Noise ($S/R=5$) 5.67 5.25 6.00 4.96 4.10 3.28
Librispeech-Noise ($S/R=0$) 15.23 13.54 14.44 13.27 11.66 9.04
Librispeech-Noise ($S/R=-5$) 49.09 47.05 47.35 46.98 43.62 33.19
Whisper GER GFD-ASR-ZH Oracle
ZH or HK dataset Greedy 5-beams (Baseline) $\text{O}_{\text{nb}}$ $\text{O}_{\text{cp}}$
Fleurs-HK (Cantonese) 7.49 6.88 7.33 6.23 5.58 4.58
NTUML2021 11.11 9.97 9.87 8.83 8.88 4.54
NTUML2021 FormosaSpeech
Full English
Whisper (short) 9.97 15.21 -
SICL (short) [49] 9.5 - -
GFD-ASR-ZH (short) 8.83 13.51 -
Whisper (long) 8.40 11.56 27.41
GFD-ASR-ZH (long) 8.18 11.43 26.72
Table 2: Comparison results on long-form and short-form.
4.3 Long-form speech recognition
As LLMs are capable in attending to long context of sequence and possess multilingual capabilities, we posit that GFD-ASR-ZH can improve on long-form audio transcription. We therefore evaluate the long-form transcription performance on NTUML2021 and FormosaSpeech. In constrast to the short-form evaluations in Table 1 in NTUML2021, we prepend all previous transcription from contiguous clips in the same lecture as prompts for both ASR and LLM to realize long form transcriptions, but truncate it when necessary for whisper due to context length limitations. In the setting of evaluations on contiguous short clips, GFD-ASR-ZH (long) consistently outperforms Whisper (long) in both NTUML2021 and FormosaSpeech, demonstrating that the long-context capability of LLM can be effectively utilized through GFD.
4.4 Context-aware speech recognition
Table 3: Results on context-aware ASR task. We set num beams=5 across all our experiments.
Method Prompt ATCO2 Librispeech FormosaSpeech Medical
ASR LLM Norm Raw Clean Other
Whisper No - 47.70 66.44 2.28 4.97 27.42 7.22
Whisper Yes - 31.34 42.37 - - - 6.24
Clairaudience[35] Yes - 28.77 - - - - 6.54
GFD-ASR-EN No Yes 38.75 52.24 2.20 4.61 26.29 6.62
GFD-ASR-EN Yes Yes 25.79 32.46 - - - 6.26
In context-aware speech recognition, we aim to investigate the capability of the fusion model in leveraging contextual information, which is highly desired in real-life contexts, where the incoming speech is expected to be bounded by domain, contain certain specific words, or follow a certain transcribing format. We utilize prompting with ASR model and the LLM model (in the GFD setting) to condition on these external hints, in three following settings.
Domain and Rare Word Prompting: We tested domain-conditioned ASR on the Medical dataset, where the symptom tags are provided along with the speech content. Results show that our GFD method with LLM prompting improves on GFD without prompting, showing prompt sensitiveness of the LLM in the GFD system. However, we did not find further improvement upon whisper prompting, compared to double-prompted GFD. For verifying rare word prompting capabilities, we used the augmented Librispeech and FormosaSpeech dataset, where a target rare word is mixed with 100 other distractors for each data point [28]. For the FormosaSpeech Dataset, we created the rare words with ChatGPT, and generate distractors with a similar approach using the training set. At this scale of distractors, it is unrealistic to prompt on Whisper, as the remaining context length is often insufficient for ASR decoding. By prompting on the LLM using GFD, we demonstrated up to 6% WERR over non-prompting methods on the Librispeech dataset, and 1.6% WERR on the FormosaSpeech Dataset.
Instruction Aware Prompting: We evaluated formatted speech recognition on the ATCO2 dataset, a dataset on air traffic control communications, which has strict regulations on call signs and transcribe formats. By incorporating a LLM, we are able to prompt it with over 4000 words of guidelines from an enitre instruction manual [^3] https://www.faa.gov/air_traffic/publications/atpubs/aim_html/chap4_section_2.html. a possibility not present with Whisper. For Whisper prompting, we include all special call signs and three example sentences extracted from the manual, where we leave details in the Appendix. Results in Table 3shows that GFD with ASR and LLM prompting obtains best results with a WERR of 17.7% compared with Whisper in the normalized (ATCO2-Norm) setting, even outperforming Clairaudience[35], a fine-tuned prompt conditioning model. We also reported scores without word conversion normalization (ATCO2-Raw), from observations that standard normalization can excessively correct errors that conflicts with the transcribing guidelines. In this setting, our gains are even larger, showing instruction following abilites of the GFD system.
4.5 Optical Character Recognition
Table 4: Evaluation on OCR task
CER $\downarrow$ Exact Match $\uparrow$
TrOCR 12.02 24.14
GFD-OCR 10.55 33.33
We demonstrate that GFD can be applied to fusing end-to-end OCR model with LLMs, using TrOCR and Mistral [32, 26]. We experiment the fusion model on a dataset extracted from National Archive Forms dataset [13], denoted NAF-Long. The NAF-Long dataset consists of images with text sequences having character lengths between 100 and 200, equivalent to around 25 to 50 text tokens. In Table 4, we show that fusing LLM to the OCR model significantly improves the OCR results by 16.7 % in character error rate reduction and 38.07 % in exact match improvement.
5 Discussion
Further comparisons with GER. The GFD algorithm relates to GER in that both algorithms perform the selection of output sequences with beam decoding. However, they differ in compute execution and the diversity of sample sequences. First, in GFD, the LLM works in parallel with the ASR decoder, and thus the computation of per step inference can be executed asynchronously with a load bounded by $\mathcal{O}(Z)+\mathcal{O}(k\cdot\max(S_{ASR},{S_{LLM}}))$, where $Z$ denotes the size of speech encoding, $k$ is the beam size, and $S_{ASR}$ and $S_{LLM}$ are ASR and LLM decoding lengths, respectively. The LLM decoding complexity expression matches that of Section 3.2, treating the vocabulary size as a constant in this dicussion. In contrast, GER operates sequentially, requiring the completion of beam decoding with ASR prior to a correction with LLM. This means the execution time of GER is bounded by $\mathcal{O}(Z)+\mathcal{O}((k+1)\cdot S_{LLM})+\mathcal{O}(k\cdot S_{ASR})$, where the additional $O(S_{LLM})$ comes from LLM decoding. Secondly, in GFD, the searched token space attended by the LLM is at least n_beams * sequence length, whereas in GER, the top-k cutoff of the ASR step does not promote diversity between the candidates, which limits the LLM search space. Aside from these differences, GFD and GER are methodologically orthogonal, allowing for their combination in the pursuit of further improvement.
Limitations. Despite promsing results, we identified three cases where GFD is most susceptible to failure, which is an exciting future direction. We categorize them into ASR errors and LLM errors.
ASR proposed errors: When ASR proposes a candidate with high probability that deviates from the phonetic constraint, the LLM may falsely pickup the sequence, inducing an error. There are two main types - semantically activated tokens and time delayed activated tokens. Semantically activated tokens are tokens that are encoded in similar output embedding space due to semantic similarity. We found that these errors are much more common in Chinese and identified some pairs from a top-k analysis of a decoding sequence, e.g., {CJK}UTF8bsmi’剛’ and’才’, ’做(什麼)’ and ’幹(什麼)’ . Time delay activated tokens are proposed tokens that are targets at a later step, where selecting them effectively skips some intermediate tokens. We observe that these tokens are much more likely to be present at the start of the sequence.
LLM probability estimation errors: LLM probability estimates are generally aligned with the logical coherence of a sequence. However, a major discrepancy arises with repeating sequences. Due to the in-context learning abilities of LLMs, they tend to significantly overestimate the likelihood of ever-repeating sequences. This could lead to mode collapse during the entire decoding process.
6 Conclusion
The Generative Fusion Decoding (GFD) framework represents a leap forward in fusing auto-regressive text models from heterogeneous token spaces. Our experiments confirm GFD’s effectiveness, particularly in ASR tasks. For tonal languages like Chinese, it achieves a mixed error rate of 8.83% on the NTUML2021 dataset, surpassing previous benchmarks. In context-aware recognition, GFD demonstrated its prowess by a 17.7% Word Error Rate Reduction (WERR) on the ATCO2 dataset when utilizing prompts to both ASR and LLM . For OCR tasks, GFD’s fusion with LLMs resulted in a notable 38.06% improvement in exact matches on the NAF-Long dataset, underscoring the framework’s ability to enhance recognition of long text sequences. These experimental outcomes highlight GFD’s robustness and adaptability, showcasing its generality to leverage the extensive training and contextual understanding of LLMs to improve text recognition across various domains and challenges. We hope that our framework will motivate further research into more fusion techniques that better leverages the existing capabilties of the pre-trained models.

### v4

@@
-Let’s Fuse Step by Step: A Generative Fusion Decoding Algorithm with LLMs for Multi-modal Text Recognition
+Let’s Fuse Step by Step: A Generative Fusion Decoding Algorithm with LLMs for Robust and Instruction-Aware ASR and OCR
 Abstract
-We introduce “Generative Fusion Decoding” (GFD), a novel shallow fusion framework, utilized to integrate Large Language Models (LLMs) into multi-modal text recognition systems such as automatic speech recognition (ASR) and optical character recognition (OCR).
-We derive the formulas necessary to enable GFD to operate across mismatched token spaces of different models by mapping text token space to byte token space, enabling seamless fusion during the decoding process.
-The framework is plug-and-play, compatible with various auto-regressive models, and does not require re-training for feature alignment, thus overcoming limitations of previous fusion techniques.
-We highlight three main advantages of GFD: First, by simplifying the complexity of aligning different model sample spaces, GFD allows LLMs to correct errors in tandem with the recognition model, reducing computation latencies.
-Second, the in-context learning ability of LLMs is fully capitalized by GFD, increasing robustness in long-form speech recognition and instruction aware speech recognition.
-Third, GFD enables fusing recognition models deficient in Chinese text recognition with LLMs extensively trained on Chinese.
-Our evaluation demonstrates that GFD significantly improves performance in ASR and OCR tasks, with ASR reaching state-of-the-art in the NTUML2021 benchmark.
-GFD provides a significant step forward in model integration, offering a unified solution that could be widely applicable to leveraging existing pre-trained models through step by step fusion.
+We propose “Generative Fusion Decoding” (GFD), a novel shallow fusion framework designed to integrate large language models (LLMs) into cross-modal text recognition systems for automatic speech recognition (ASR) and optical character recognition (OCR).
+We derive the necessary formulations to enable GFD to operate across mismatched token spaces of different models by calculating likelihood at the byte level, thereby enabling seamless fusion and synchronous progression during the decoding process.
+GFD is plug-and-play by design, making it readily compatible with various auto-regressive models without the need for any re-training.
+GFD proves effective for general ASR and OCR tasks through intermediate and frequent interactions with LLMs, surpassing cascaded methods in English and Mandarin benchmarks.
+In addition, GFD transfers in-context learning abilities of LLMs and allows for adaptive ASR in instruction-aware and long-context settings, yielding significant WER reductions of up to 17.7%.
 [^1] Code is available at https://github.com/mtkresearch/generative-fusion-decoding
+Let’s Fuse Step by Step: A Generative Fusion Decoding Algorithm with LLMs for Robust and Instruction-Aware ASR and OCR
+Chan-Jan Hsu∗1  Yi-Chang Chen∗1  Feng-Ting Liao1 Pei-Chen Ho2Yu-Hsiang Wang2Po-Chun Hsu1Da-shan Shiu1 Equal contribution  1MediaTek Research  2Internship at MediaTek Research {chan.hsu, yi-chang.chen, ft.liao, pochun.hsu, ds.shiu}@mtkresearch.com
 1 Introduction
-In recent years, integrating Large Language Models (LLMs) into multi-modal systems has emerged as a frontier, offering substantial advancements in applications such as automatic speech recognition (ASR) [40], visual question answering (VQA) [36], reinforcement learning [50], and generalist agent [42].
-The robust language model capabilities that the LLM offers make it an ideal candidate for integration with other models, many of which are “undertrained” in the textual domain, compared to trillions of tokens consumed by the LLMs.
-Obtaining sufficient quality paired data and training resources in a huge endeavor for text recognition systems such as automatic speech recognition (ASR) and optical character recognition (OCR).
-The integration of heterogeneous models has been explored extensively in ASR literature, as exemplified by strategies such as shallow fusion, late fusion, mid fusion, and early fusion.
-Works in shallow fusion [9, 27, 10] interpolates text recognition outputs with LM outputs at each step of beam search.
-Late fusion approaches typically engage LLMs to score the initial sentence generated by the recognition model.
-The scores are then utilized for fusing with recognition probabilities for joint transcription [6], for re-ranking the the constrained set of hypothesis as work in re-scoring [46], for generating new hypothesis options through Generative Error Correction (GER) [5].
-Notably, methods in mid-fusion propose fusing models through model attentions to integrate LLM into the decoding process of the recognition model [41, 37], and work in early-fusion has proposed concatenating speech embeddings or image embeddings along with LLM embedings as inputs to the LLM for ASR and VQA [17, 7].
-However, the aforementioned fusion methods bear several challenges.
-The early fusion techniques are often limited by a low correction rate due to the restricted range of available choices for re-evaluation.
-Mid-fusion discards the whisper decoder, which has been shown to have denoising ability[19].
-Additionally, late fusion with GER may lead to hallucination due to the LLM as the LLM generation is devoid of auditory inputs.
-Due to the requirements for shallow and late fusion to align output spaces, and the need for mid and early fusions to involve joint training with additional paired data, prevent an LLM from being used as is with existing text-recognition models without proper feature alignment through trainings.
 Figure 1: The GFD integrated framework.
-The framework aims to integrate pre-trained text-recognition models with LLMs to augment the recognition capabilities.
-However, direct fusion is impeded by the discrepancy in token spaces between the two model types.
-To overcome this obstacle, GFD is introduced, which converts the token space into a byte space, thereby facilitating incremental fusion.
-Equations 7 and 8 are employed to compute the fusion probability, enabling a stepwise fusion process.
-The symbols $\mathcal{Z}^{\texttt{TR}}$ and $\mathcal{Z}^{\texttt{LM}}$ denote the contextual information pertinent to the text-recognition models and the language models, respectively.
-Notably, $\mathcal{Z}^{\texttt{TR}}$ may correspond to either an audio input for ASR or a text image for OCR, whereas $\mathcal{Z}^{\texttt{LM}}$ typically represents a textual prompt.
-To address the aforementioned challenges, we introduce a novel shallow fusion framework, termed “Generative Fusion Decoding” (GFD).
-GFD resolves the token space mismatch between models by bridging heterogeneous spaces through the expected score values of plausible sequences and mapping text token space to byte token space.
-This allows GFD to seamlessly fuse LLMs with text recognition models during the decoding process.
-Our unified framework is plug-and-play, applicable to text recognition tasks such as ASR and OCR, and compatible with diverse auto-regressive text recognition models as well as language models.
-We highlight three main advantages of GFD.
-First, GFD addresses key limitation of previous fusion techniques — namely the complexity of aligning different model sample spaces via re-training associated with prior deep integration [41] or two-pass [46] strategies.
-It allows LLMs to participate in the sampling process as conventional shallow fusion, providing real-time correction of errors generated by the text recognition model, thereby extending the influence of LLMs beyond the complete outputs of the recognition model.
-Secondly, as modern LLMs are trained on trillions of multilingual text tokens [22, 26], generally around 3 orders of magnitude more than end-to-end ASR or OCR models [40], LLMs in GFD can be seen as a strong scoring model in the step by step decoding that improves the multi-lingual outputs of ASR and OCR systems.
-Lastly, because of the in-context learning ability of LLMs, the GFD fusion system with LLMs and ASR can readily leverage domain information as prompts effectively for domain adaptation similarly to prior work through re-scoring [14] or instruction tuning [35].
-To evaluate the efficacy of our algorithm, we implemented the GFD for ASR and OCR tasks.
-In the ASR task, we evaluated on the conventional ASR task as shown in Section 4.2 for English and Chinese audios.
-The evaluation of Chinese ASR is motivated by existing literature [40], which indicates a performance outlier in Chinese compared to other languages despite abundant paired-training data.
-We hypothesized that LLMs could enhance Chinese ASR by offering additional scoring to differentiate between semantically distinct but homophonous words.
-Furthermore, we demonstrate that ASR with GFD can improve recognition for long-form audio through leveraging the long-context ability of LLMs (Section 4.3); for context-aware recognition for rare words and domain tags via in-context learning capabilities of LLMs (Section 4.4).
-Lastly, application to the OCR task, particularly targeting long text OCR for historical documents, demonstrates that OCR with GFD can significantly improve recognition accuracy (Section 4.5).
+The framework aims to integrate pre-trained text-recognition models (ASR/OCR) with LLMs to augment the recognition capabilities.
+A key challenge in this integration lies in the mismatch between the token spaces of the two model types, which prevents direct fusion.
+To address this, we derive a formulation (Section 3.1, Equation 7, Equation 8) that enables GFD to compute likelihoods at the byte level, allowing effective fusion during the decoding stage.
+Here, $\mathcal{Z}^{\texttt{TR}}$ and $\mathcal{Z}^{\texttt{LM}}$ denote the contextual information from the text-recognition model and the LLM, respectively.
+Integrating large language models (LLMs) into multi-modal systems has recently emerged as a frontier, significantly advancing applications such as automatic speech recognition (ASR) Radford et al.
+(2023), visual question answering (VQA) Liu et al.
+(2023), and reinforcement learning Yang et al.
+(2023d).
+Despite their robust capabilities, integrating LLMs with text recognition systems like ASR and OCR poses challenges due to the need for high-quality paired data and extensive training resources.
+Modern LLMs are trained on trillions of text tokens Hsu et al.
+(2024); Jiang et al.
+(2023), far exceeding the data used for end-to-end ASR or OCR models Radford et al.
+(2023).
+Various fusion strategies have been explored in ASR literature, including shallow fusion Chen et al.
+(2023b); Kannan et al.
+(2018); Choudhury et al.
+(2022), late fusion Chen et al.
+(2024b, a); Xu et al.
+(2022), mid fusion Radhakrishnan et al.
+(2023); Liu et al.
+(2024), and early fusion Fathullah et al.
+(2024); Chen et al.
+(2023a).
+However, these methods face challenges such as discarding the ASR decoder’s denoising abilities Gong et al.
+(2023) and requiring aligned token spaces.
+Volatility of model from further training is also a concern when dealing with extensively trained models.
+To address these challenges, we introduce a novel shallow fusion framework called “Generative Fusion Decoding” (GFD).
+GFD operates across mismatched token spaces by calculating likelihood at the byte level, enabling seamless integration of LLMs with text recognition models during the synchronous decoding process (Section 3.1).
+This plug-and-play framework allows LLMs to correct text recognition errors in real-time (Section 3.2), broadening the exploration space and improving recognition accuracy.
+Empirically, GFD is effective on general ASR, especially in challenging scenarios like homophones in Mandarin and code-switching Yang et al.
+(2023c) (Section 4.2).
+In addition, GFD transfers long-context awareness and in-context learning Brown et al.
+(2020b) of LLMs and allows for adaptive ASR.
+GFD maintains semantic consistency in long-form audio by leveraging transcription history for contextual biasing (Section 4.3).
+By using controlling prompts such as domain tags, rare words, and explicit instructions, domain sensitivity and instruction awareness is exhibited across various benchmarks (Section 4.4).
+To the best of our knowledge, this unique aspect of LLM integration has not been reported in prior work Chen et al.
+(2024b); Hu et al.
+(2024); Mittal et al.
+(2024); Hori et al.
+(2025).
+Further comparisons with existing methods are discussed in Section 5.
+The contributions of this work are summarized as follows:
+  • We derive a novel algorithm – GFD, which enables intermediate LLM interaction during the decoding process in text recognition.
+  • GFD improves performance on various ASR scenarios and OCR, which is orthogonal to improvements from previous approaches.
+  • The robustness of GFD is demonstrated in long-context and instruction-based ASR tasks, which fully utilizes long-range semantic awareness of LLMs.
+To the best of our knowledge, this has not been reported in prior work with LLM integration.
+  • We provide detailed analysis on the performance and the time efficiency of GFD.
 2 Related work
 2.1 Model fusion
-Training a multi-objective model from scratch is often costly [3, 2, 1, 15].
+Training a multi-objective model from scratch is often costly Bapna et al.
+(2021, 2022); Alayrac et al.
+(2022); Driess et al.
+(2023).
 Consequently, researchers have pivoted towards combining existing models with different modalities to improve accuracy without the prohibitive costs of building new systems from the ground up.
@@
 The technique have evolved significantly in recent years, encompassing a variety of approaches designed to integrate different models to enhance performance.
-Shallow fusion or late fusion, often employed in ASR, combines end-to-end ASR models with external language models at the decoding level, improving recognition accuracy without altering the underlying ASR architecture [27, 25, 6, 52].
-However, due to the heterogeneous sample spaces of models, the prerequisite of shallow and late fusion requires aligning sample spaces of model distributions, enabled through fine-tuning a projection module [6].
-Late fusion training methods may suffer from modality laziness problem in tasks where uni-modal priors are meaningful [16].
-Work in re-scoring N-best hypothesis from the first pass inference with an external language model has been shown effective in reducing recognition errors [43, 23, 46].
-However, the method prunes the correct prefix early on away from N-best list and requires a second pass in the decoding process.
-Alternatively, deep fusion integrates models at the level of hidden feature, requiring fine-tuning models to fuse deep features [20].
-Cross-modal fusion, similar to deep fusion, integrates pre-trained end-to-end ASR model with LLM [41, 51, 33] or vision model with LLM [7, 36] via learning a joint representation with large amount of extra paired audio-text or image-text data.
-Our newly proposed approach—GFD—decodes in homogeneous sequence element space with byte strings, and therefore relaxes the alignment of sample spaces of distributions without any model training.
-Without pre-training or fine-tuning, we ensure that the individual pre-trained model performance is guaranteed on their respective task, and is not influenced by volatility of further training.
+Deep fusion integrates models at the level of hidden features, requiring fine-tuning models to fuse deep features Gulcehre et al.
+(2015).
+Cross-modal fusion, similar to deep fusion, integrates pre-trained end-to-end ASR model with LLM Radhakrishnan et al.
+(2023); Yu et al.
+(2023); Li et al.
+(2023b) or vision model with LLM Chen et al.
+(2023a); Liu et al.
+(2023) via learning a joint representation with large amount of extra paired audio-text or image-text data.
+In contrast, shallow fusion or late fusion, often employed in ASR, combines end-to-end ASR models with external language models at the decoding level, improving recognition accuracy without altering the underlying ASR architecture Kannan et al.
+(2018); Huang et al.
+(2024); Chen et al.
+(2024b); Zhang et al.
+(2023).
+However, due to the heterogeneous sample spaces of models, the prerequisite of shallow and late fusion requires aligning sample spaces of model distributions, enabled through fine-tuning a projection module Chen et al.
+(2024b).
+Late fusion training methods may suffer from modality laziness problem in tasks where uni-modal priors are meaningful Du et al.
+(2023).
+Concurrently with our work, step-by-step synchrounous late fusion methods are explored Mittal et al.
+(2024); Hori et al.
+(2025).
+Departing from these efforts, which constrain scoring to specific decoding configurations, our approach addresses the problem from the byte sequence perspective, generalizing the rescoring process to support arbitrary input sequences.
+Another line of research integrates LLMs in a cascaded fashion, where the LLM rescores or rewrites based on the N-best hypotheses generated by the first-pass ASR model.
+While this approach has proven effective in reducing recognition errors Sainath et al.
+(2019); Hu et al.
+(2020); Xu et al.
+(2022), it is limited by the inherently low representation capacity of the N-best list and introduces additional computational latency from the second-pass decoding.
+Our newly proposed approach, GFD, operates in the space of homogeneous sequence elements, thereby removing the need for strict token-level alignment.
+Further, by keeping the model architecture intact, including tokenizers and embedding, we ensure that each individual pre-trained model’s performance on its respective task is preserved and not affected by the instability that may arise from additional training.
+This property becomes increasingly critical when integrating with large language models (LLMs), which are typically trained on trillions of tokens using carefully refined data curricula and annealed learning rates Dubey et al.
+(2024).
 2.2 Contextual conditioning
-Auto-regressive LLMs have exhibited capabilities in in-context learning [39], instruction following [38], and knowledge synthesis [44].
-Such capabilities have been applied to solving domain adaptation in speech recognition for rare words or out-of-domain context through contextual biasing [10] and prompting fine-tuned models [35, 47, 33, 47].
-With GFD, such problem setting can be tackled by leveraging a performant LLM as is through prompting without further fine-tuning.
-2.3 Chinese ASR
-One of the most significant challenges in developing ASR systems for the Chinese language stems from its highly homophonous nature [30, 29, 8].
-Unlike English, where there is a larger variety of phonemes and a relatively consistent correspondence between spelling and sound, Chinese relies on a limited set of tones and syllables to represent thousands of characters.
-Consequently, Chinese ASR systems must not only accurately capture the tonal nuances but also analyze the linguistic context to disambiguate these homophones.
-This requires sophisticated LMs that can effectively leverage contextual clues to predict the intended characters.
-The integration of LLMs has shown promise in addressing these challenges [11, 31, 34].
-GFD is also expected to enhance the performance of Chinese ASR by utilizing the contextual conditioning capabilities of LLMs.
+Auto-regressive LLMs have exhibited capabilities in in-context learning Radford et al.
+, instruction following Ouyang et al.
+(2022), and knowledge synthesis Liu et al.
+(2020).
+Such capabilities have been applied to solving domain adaptation in speech recognition for rare words or out-of-domain context through contextual biasing Choudhury et al.
+(2022) and prompting fine-tuned models Liao et al.
+(2023); Yang et al.
+(2023a); Li et al.
+(2023b); Yang et al.
+(2023a).
+Using GFD, this problem can be addressed by directly leveraging a high-performing LLM through prompting, without the need for additional fine-tuning.
+2.3 Mandarin ASR
+One of the most significant challenges in developing ASR systems for Mandarin stems from its highly homophonous nature Lee and Chen (1997); Lee (2003); Chen et al.
+(2022).
+Unlike English, where there is a larger variety of phonemes and a relatively consistent correspondence between spelling and sound, Mandarin relies on a limited set of tones and syllables to represent thousands of characters.
+Consequently, Mandarin ASR systems must not only accurately capture the tonal nuances but also analyze the linguistic context to disambiguate these homophones.
+The integration of LLMs has shown promise in addressing these challenges Chung et al.
+(2023); Leng et al.
+(2023); Li et al.
+(2024), and GFD adopts the same ideology by leveraging the contextual conditioning capabilities of LLMs to enhance Chinese ASR performance.
 3 Method
@@
 $\{T_{s}\}^{*}=\underset{\{T_{s}\}}{\mathrm{arg\,max}}\log\mathbb{P}(\{T_{s}\},\mathcal{Z}),$ (1)
-where $\{T_{s}\}$ represents the sequence of tokens generated by the model, and $\mathcal{Z}$ represents the given context or conditioning information, such as audio for speech recognition models [40], images for vision-language-models [1], and prompts for typical language models [4].
+where $\{T_{s}\}$ represents the sequence of tokens generated by the model, and $\mathcal{Z}$ represents the given context or conditioning information, such as audio for speech recognition models Radford et al.
+(2023), images for vision-language-models Alayrac et al.
+(2022), and prompts for typical language models Brown et al.
+(2020a).
 Auto-regressive generation is one approach to realize conditional text generation.
 In this approach, the auto-regressive model is conditioned on the previously generated tokens to generate the next token sequentially.
-Therefore, the probability $\log\mathbb{P}(\{T_{s}\}|\mathcal{Z})$ is typically decomposed using the chain rule of probability, as follows:
+Therefore, the probability $\log\mathbb{P}(\{T_{s}\}|\mathcal{Z})$ is typically decomposed using the chain rule of probability as follows:
 $\log\mathbb{P}(\{T_{s}\},\mathcal{Z})=\sum_{s=1}^{S}\log\mathbb{P}(T_{s}\mid T_{<s},\mathcal{Z}),$ (2)
 where $T_{s}$ is the token at position $s$ in the sequence, and $T_{<s}$ represents all the tokens preceding position $s$.
-In real-world applications, it is impracticable to enumerate all possible token sequences, so beam search is typically used as an approximation method.
-In the setting of shallow fusion, we attempt to let multiple models jointly determine the sequence, as shown in the following formula rewrote from Equation (1):
-$\{T_{s}^{\texttt{fuse}}\}^{*}$ $=\underset{\{T_{s}^{\texttt{fuse}}\}}{\mathrm{arg\,max}}\sum_{m}\lambda_{m}\log\mathbb{P}_{m}(\{T_{s}^{(m)}\}=\{T_{s}^{\texttt{fuse}}\},\mathcal{Z}^{(m)})$ (3)
-where $\{T_{s}^{\texttt{fuse}}\}$ represents the fused sequence of tokens generated by combining the outputs of multiple models, $\lambda_{m}$ is a weighting factor for the $m$-th model, $p_{m}$ denotes the probability distribution of the $m$-th model and $\mathcal{Z}^{(m)}$ represents the context or conditioning information specific to the $m$-th model.
-When the sample spaces of models are the same, Equations (3) and (2) can be combined to realize incremental fusion.
+In real-world applications, it is impracticable to enumerate all possible token sequences, so beam search is typically employed as an approximate strategy to efficiently explore the most likely sequences without exhaustive computation.
+In the setting of shallow fusion, multiple models are combined to jointly determine the sequence, as expressed in the following formula, which is a reformulation of Equation (1):
+$\{T_{s}^{\texttt{fuse}}\}^{*}$ (3)
+$=\underset{\{T_{s}^{\texttt{fuse}}\}}{\mathrm{arg\,max}}\sum_{m}\lambda_{m}\log\mathbb{P}_{m}(\{T_{s}^{(m)}\}=\{T_{s}^{\texttt{fuse}}\},\mathcal{Z}^{(m)})$
+where $\{T_{s}^{\texttt{fuse}}\}$ represents the fused sequence of tokens generated by combining the outputs of multiple models, $\lambda_{m}$ is a weighting factor for the $m$-th model, $\mathbb{P}_{m}$ denotes the probability distribution of the $m$-th model and $\mathcal{Z}^{(m)}$ represents the context or conditioning information specific to the $m$-th model.
+When the sample spaces of models are the same, Equations (3) and (2) can be combined to realize incremental fusion Chen et al.
+(2024b).
 If the models have different sample spaces due to a mismatch in token spaces, there is no simple way to achieve incremental fusion.
@@
 To address these challenges, we have introduced a probability transformation, denoted as $\mathcal{M}^{(m)}$, that converts token-level representations into byte-level representations:
-$\mathcal{M}^{(m)}:\mathbb{P}_{m}(\{T_{s}\},\mathcal{Z}^{(m)})\longrightarrow\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)}),$ (4)
+$\mathcal{M}^{(m)}:\mathbb{P}_{m}(\{T^{(m)}_{s}\},\mathcal{Z}^{(m)})\longrightarrow\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)}),$ (4)
 where $\{B_{l}\}$ represents the sequence of bytes after the transformation, and $l$ denotes the position in the byte sequence.
@@
 The byte-level fusion can then be performed using a similar approach to Equation (3), but with the byte-level probabilities:
-$\{B_{l}^{\texttt{fuse}}\}^{*}$ $=\underset{\{B_{l}^{\texttt{fuse}}\}}{\mathrm{arg\,max}}\sum_{m}\lambda_{m}\log\mathbb{P}_{m}(\{B_{l}\}=\{B_{l}^{\texttt{fuse}}\},\mathcal{Z}^{(m)}).$ (5)
-To operationalize the probability transformation $\mathcal{M}^{(m)}$, we define a mapping from the token-level probabilities to the byte-level probabilities.
+$\{B_{l}^{\texttt{fuse}}\}^{*}$ (5)
+$=\underset{\{B_{l}^{\texttt{fuse}}\}}{\mathrm{arg\,max}}\sum_{m}\lambda_{m}\log\mathbb{P}_{m}(\{B_{l}\}=\{B_{l}^{\texttt{fuse}}\},\mathcal{Z}^{(m)}).$
+To realize the probability transformation $\mathcal{M}^{(m)}$, we define a mapping from the token-level probabilities to the byte-level probabilities.
 This mapping takes into account the prefix relationship between the token sequence and the byte sequence.
@@
 This relationship is formalized in the following equation:
-$\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})=\sum_{\{B_{l}\}}\biggr{[}\prod_{s}\mathbb{P}_{m}(T_{s}^{(m)}\mid T_{<s}^{(m)},\mathcal{Z}^{(m)})\biggr{]}_{\{T_{s}\}}\times\mathbb{1}\bigl{(}\{T_{s}\}\texttt{.prefix}=\{B_{l}\}\bigl{)},$ (6)
-where .prefix is a function that maps a token sequence to its corresponding byte sequence prefix, and $\mathbb{1}$ is the indicator function that is 1 if the token sequence has the specified byte sequence as a prefix and 0 otherwise.
-The indicator function ensures that only those token sequences that can be mapped to the given byte sequence contribute to the byte-level probability.
+$\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})$ (6)
+=$=\sum_{\{T^{(m)}_{s}\}}\biggr{[}\prod_{s}\mathbb{P}_{m}(T_{s}^{(m)}\mid T_{<s}^{(m)},\mathcal{Z}^{(m)})\biggr{]}_{\{T^{(m)}_{s}\}}$
+×$\times\mathbbm{1}\bigl{(}\{T^{(m)}_{s}\}\texttt{.pref}=\{B_{l}\}\texttt{~{}AND~{}}T^{(m)}_{<s}\texttt{.pref}\neq\{B_{l}\}\bigl{)},$
+where .pref is a function that checks whether a sequence $A$ has sequence $B$ as its prefix, and $\mathbbm{1}$ is the indicator function that converts the boolean value of the inner loop to integers ($true\rightarrow 1$, $false\rightarrow 0$).
+The entire indicator function with two conditions ensures that only the minimal token sequences covering the target byte sequence $\{B_{l}\}$ contribute to the byte-level probability.
 In Equation (6), the complexity remains high at $\mathcal{O}(V^{S})$, where $V$ represents the vocabulary size of tokens and $S$ is the sequence length, for identifying token sequences that match the specified criteria of the indicator.
@@
 The alternative tokens are denoted as $T^{\prime}_{i}$.
-To reduce complexity, we posit that the main token sequence and its branching alternatives, as shown in Figure 2, dominate the probability contribution, where the main token sequence is produced by applying model tokenization on the byte string and the alternative tokens are essentially look-aheads for potential main tokens that may emerge as the decoding progresses.
+This complexity can still be greatly reduced, by eliminating terms with near 0 probability.
+We posit that the main token sequence and its branching alternatives, as shown in Figure 2, dominate the probability contribution.
+The main token sequence is produced by applying model tokenization on the byte string and the alternative tokens are essentially look-aheads for potential main tokens that may emerge as the decoding progresses.
 The significance of the main token sequence is justified by its alignment with the model’s inputs during the original pretraining phase; any other slicing method is penalized in terms of probability due to its lack of representation in the training data.
-Based on this assumption, we can narrow down the search for token sequences that meet the criteria of $\{T_{s}\}\texttt{.prefix}=\{B_{l}\}$ to only the main token sequence and its branching alternative tokens.
-We define the main token sequence as $\{\hat{T}_{s}\}$.
+Based on this assumption, we can narrow down the search for token sequences that meet the criteria of ($\bigl{(}\{T^{(m)}_{s}\}\texttt{.pref}=\{B_{l}\}\texttt{~{}AND~{}}T^{(m)}_{<s}\texttt{.pref}\neq\{B_{l}\}\bigl{)}$ to only the main token sequence and its branching alternative tokens.
+We define the main token sequence as $\{\hat{T}^{(m)}_{s}\}$.
 Given this simplification, we can approximate the byte-level probability $\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})$ by considering only the main token sequence and its immediate alternatives that share the same prefix with the byte sequence $\{B_{l}\}$.
 This approximation significantly reduces the computational complexity to $\mathcal{O}(V\times S)$ and is expressed in the following equation:
-$\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})\approx$  $\texttt{~{}}\mathbb{P}_{m}(\hat{T}_{1}\mid\hat{T}_{<1},\mathcal{Z}^{(m)})\times\bigl{[}$ (7)
- $\texttt{~{}~{}~{}}\mathbb{P}_{m}(\hat{T}_{2}\mid\hat{T}_{<2},\mathcal{Z}^{(m)})\times\bigl{[}$
+$\mathbb{P}_{m,approx}(\{B_{l}\},\mathcal{Z}^{(m)})\approx$ (7)
+ $\texttt{~{}}\mathbb{P}_{m}(\hat{T}^{(m)}_{1}\mid\hat{T}^{(m)}_{<1},\mathcal{Z}^{(m)})\times\bigl{[}$
+ $\texttt{~{}~{}}\mathbb{P}_{m}(\hat{T}^{(m)}_{2}\mid\hat{T}^{(m)}_{<2},\mathcal{Z}^{(m)})\times\bigl{[}$
 […]
- $\texttt{~{}~{}~{}~{}~{}~{}~{}}\mathbb{P}_{m}(\hat{T}_{S}\mid\hat{T}_{<S},\mathcal{Z}^{(m)})$
- $\texttt{~{}~{}~{}~{}~{}~{}~{}}+\sum_{t}\mathbb{P}_{m}(t\mid\hat{T}_{<S},\mathcal{Z}^{(m)})\cdot\mathbb{1}\bigl{(}\{\hat{T}_{<S},t\}\texttt{.prefix}=\{B_{l}\}\bigl{)}$
+ $\texttt{~{}~{}~{}}\mathbb{P}_{m}(\hat{T}^{(m)}_{S}\mid\hat{T}^{(m)}_{<S},\mathcal{Z}^{(m)})$
+ $\texttt{~{}~{}~{}}+\sum_{t}\mathbb{P}_{m}(t\mid\hat{T}^{(m)}_{<S},\mathcal{Z}^{(m)})\cdot\mathbbm{1}\bigl{(}\{\hat{T}^{(m)}_{<S},t\}\texttt{.pref}=\{B_{l}\}\bigl{)}$
 […]
- $\texttt{~{}~{}~{}}\bigl{]}+\sum_{t}\mathbb{P}_{m}(t\mid\hat{T}_{<2},\mathcal{Z}^{(m)})\cdot\mathbb{1}\bigl{(}\{\hat{T}_{<2},t\}\texttt{.prefix}=\{B_{l}\}\bigl{)}$
- $\texttt{~{}}\bigl{]}+\sum_{t}\mathbb{P}_{m}(t\mid\hat{T}_{<1},\mathcal{Z}^{(m)})\cdot\mathbb{1}\bigl{(}\{\hat{T}_{<1},t\}\texttt{.prefix}=\{B_{l}\}\bigl{)},$
-where $t$ represents an alternative token that could follow the sequence $\hat{T}_{<s}$ and meet the criteria of $\{T_{s}\}\texttt{.prefix}=\{B_{l}\}$.
-Eventually, substituting Equation (7) into (5) can realize generative fusion decoding (GFD).
+ $\texttt{~{}~{}}\bigl{]}+\sum_{t}\mathbb{P}_{m}(t\mid\hat{T}^{(m)}_{<2},\mathcal{Z}^{(m)})\cdot\mathbbm{1}\bigl{(}\{\hat{T}^{(m)}_{<2},t\}\texttt{.pref}=\{B_{l}\}\bigl{)}$
+ $\texttt{~{}}\bigl{]}+\sum_{t}\mathbb{P}_{m}(t\mid\hat{T}^{(m)}_{<1},\mathcal{Z}^{(m)})\cdot\mathbbm{1}\bigl{(}\{\hat{T}^{(m)}_{<1},t\}\texttt{.pref}=\{B_{l}\}\bigl{)},$
+where $t$ represents an alternative token from the token set of the modality $m$.
+We use $\{\hat{T}^{(m)}_{<s},t\}$ to denote the concatenated sequence of $\hat{T}^{(m)}_{<s}$ and $t$, and this sequence must meet the criteria of leading with $\{B_{l}\}$ to be considered.
+Eventually, by substituting Equation (7) into $\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})$ of (5), we have successfully realized generative fusion decoding (GFD).
+We show that this function is incrementally calculable in Appendix A.1, and thus yields the same time complexity as standard LLM rescoring without branching.
 In summary, our proposed method for conditional text generation, GFD, through late fusion and byte-level probability transformation offers a novel way to integrate the outputs of multiple models with different token spaces.
@@
 This is achieved by fusing pre-trained text-recognition models with LLMs to enhance recognition capability.
-In such scenarios, the proposed tokens of ASR and OCR model are conditioned on a modal signal $\mathcal{Z}^{\texttt{TR}}$ while that of the LLM is not.
-Therefore, the same next-byte probability estimation from the LLM and text recognition model might stem from different token choices.
-Although the difference would converge eventually as the sequence rolls out, given limited beams in the beam search process, the potential overestimation of implausible words due to scores contributed by LLM spikes the probabilities of common word-leading bytes such as ’t’ and ’r’, and produces locally unstable rankings that may result in false eliminations of correct beams in decoding steps.
-In response to this challenge, we have adopted text recognition models (ASR and OCR) to propose the next byte/token, while employing LLMs to provide delayed corrective feedback on past tokens.
-The fusion decoding methodology we have adopted, derived from Equation (5), is delineated in the following formula:
-$\{B_{1},..,B_{t}\}^{*}=\underset{\{B_{1},..,B_{t}\}}{\mathrm{arg\,max}}\bigl{[}(1-r)\cdot\log\mathbb{P}^{\texttt{TR}}(\{B_{1},..,B_{t}\},\mathcal{Z}^{\texttt{TR}})$ (8)
+Essentially, the text recognition models (ASR and OCR) propose sequences for the LLMs to provide scoring feedback.
+To limit the number of proposals scored by the LLM for reasonable time complexity, we introduce a delayed corrective feedback loop to coordinate the two models, characterized by a dynamic shifting value $k$.
+Based on Equation (5), the fusion decoding methodology used in our experiments is given by the following formula:
+$\{B_{1},..,B_{t}\}^{*}$ (8)
+$=\underset{\{B_{1},..,B_{t}\}}{\mathrm{arg\,max}}\bigl{[}$ $(1-r)\cdot\log\mathbb{P}^{\texttt{TR}}(\{B_{1},..,B_{t}\},\mathcal{Z}^{\texttt{TR}})$
 $+r\cdot\log\mathbb{P}^{\texttt{LM}}(\{B_{1},..,B_{t-k}\},\mathcal{Z}^{\texttt{LM}})\bigl{]},$
-where $\{B_{1},\ldots,B_{t}\}^{*}$ represents the optimal sequence of bytes up to and including position $t$, $p^{\texttt{TR}}$ and $p^{\texttt{LM}}$ are the probability distributions of the text-recognition models and the language models, respectively, $\mathcal{Z}^{\texttt{TR}}$ and $\mathcal{Z}^{\texttt{LM}}$ are the contextual information for the text-recognition and language models, respectively, $r$ is a weighting factor that balances the influence of the text-recognition models and the language models, and $k$ depends on the length of the byte/token proposal.
-As shown in Figure 1, $\mathcal{Z}^{\texttt{TR}}$ can be either an audio or a text image, and $\mathcal{Z}^{\texttt{LM}}$ can be a prompt.
-4 Experiment
+where $\{B_{1},\ldots,B_{t}\}^{*}$ represents the optimal sequence of bytes up to and including position $t$, $p^{\texttt{TR}}$ and $p^{\texttt{LM}}$ are the probability distributions of the text-recognition models and the language models, respectively, $\mathcal{Z}^{\texttt{TR}}$ and $\mathcal{Z}^{\texttt{LM}}$ are the contextual information for the text-recognition and language models, respectively, $r$ is a weighting factor that balances the influence of the text-recognition models and the language models, which is determined via grid search on a small scale experiment (Appendix  A.2).
+$k$ is optimally selected to be equal to the length of the last token of the proposal from $\mathcal{M}^{\texttt{TR}}$ (Appendix  A.3).
+We note that even with a shifting value $k$, the necessity of equation 7 still holds.
+4 Experiments
+Whisper Re-ranking GER GFD-ASR-EN Oracle
+EN dataset Greedy 5-beams $\text{O}_{\text{nb}}$ $\text{O}_{\text{cp}}$
+Librispeech-Clean 2.30 2.28 2.20 2.41 2.29 1.91 1.63
+Librispeech-Other 5.23 4.97 4.86 5.30 4.99 4.20 3.43
+Medical 7.30 7.22 7.22 7.49 6.74 6.17 5.05
+Librispeech-Noise ($S/R=10$) 3.50 3.12 3.27 3.14 3.02 2.32 1.94
+Librispeech-Noise ($S/R=5$) 5.67 5.25 5.40 5.38 4.96 4.10 3.28
+Librispeech-Noise ($S/R=0$) 15.23 13.54 13.69 13.70 13.27 11.66 9.04
+Librispeech-Noise ($S/R=-5$) 49.09 47.05 47.04 47.35 46.98 43.62 33.19
+Whisper Re-ranking GER GFD-ASR-ZH Oracle
+ZH or HK dataset Greedy 5-beams $\text{O}_{\text{nb}}$ $\text{O}_{\text{cp}}$
+Fleurs-HK (Cantonese) 7.49 6.88 7.00 7.33 6.23 5.58 4.58
+NTUML2021 11.11 9.97 9.68 9.87 8.83 8.88 4.54
+Table 1: Performance for short-form speech recognition.
+The table presents Word Error Rate (WER) for EN and Mixed Error Rate (MER) for ZH/HK.
+Bold values indicate the best performance excluding the Oracle column.
+Re-ranking and GER utilize LLMs with 5-beam outputs from Whisper, whereas GFD integrates LLMs during the decoding process.
+Librispeech-Noise (S/R=0)
+Whisper-5beams 13.54
+RobustGER Hu et al. (2024) 13.20
+GFD 13.27
+RobustGER+GFD 13.03
+Table 2: Comparisons on GER and GFD. GFD and GER improvements similarly on Whisper, where ensembling of both approaches performs best.
 4.1 Experimental setup
 We evaluate the application of GFD to ASR and OCR tasks.
-For ASR task, we benchmark datasets in English, Chinese, and Cantonese.
-The deliberate choice of Chinese and Cantonese is due to that the languages are homophonous and tonal, challenging for ASR systems.
-This complexity is corroborated by the Fleurs experiment in Whisper [40], where the Chinese word error rate is well above the regressed word error rate in comparison to evaluations in other language at the same amount of pre-training data.
-Of all tested languages in Whisper, it is the only large-scale language (> 10k hours audio) with such a phenomenon.
+For ASR task, we benchmark datasets in English, Taiwanese Mandarin, and Cantonese.
+The deliberate selection of Taiwanese Mandarin and Cantonese is due to their homophonic and tonal characteristics, which reveal robustness shortcomings of ASR systems.
+This complexity is corroborated by the Fleurs experiment in Whisper Radford et al.
+(2023), where the Chinese word error rate is well above the regressed word error rate in comparison to evaluations in other language at the same amount of pre-training data.
+Of all tested languages in Whisper, it is the only large-scale language (more than 10k hours audio) with such a phenomenon.
 For OCR, we benchmark image dataset containing long sequence of text as we hypothesize that LLM provide semantic information to an OCR model of recognizing long text sequences.
-In terms of the evaluated models, we choose Whisper-large-v2 as the ASR model for greedy and beam search.
-For our proposed GFD, we use Mistral [26] for English datasets and Breeze [22] for Chinese and Cantonese datasets.
-We term the GFD fused model with Whisper and Mistral as GFD-ASR-EN; that with Whisper and Mistral as GFD-ASR-ZH.
-In addition, we benchmark GER, based on task-activating prompting method [5, 48], with Mistral-Instruct and Breeze-Instruct models without finetuning.
-We include two oracle word error rates following previous work[24], where the N-Best Oracle $o_{nb}$ denotes the error rate calculated with the best candidate in the N-best list, and the Compositional Oracle $o_{cp}$ is the best achievable word error rate using all tokens N-best the list.
-As to OCR, we utilize TrOCR [32], a pre-trained OCR model, and denote the fused model GFD-OCR.
-We benchmark models on various datasets, including Librispeech [21], Medical [18], ATCO2 [45], Fleurs [12], NTUML2021[49], and FormosaSpeech  [^2] https://huggingface.co/datasets/MediaTek-Research/formosaspeech for ASR; NAF [13] for OCR.
-Librispeech is a collection of corpus from audiobooks with subsets Clean and Other.
-The Librispeech-Noise is its noised variant, featuring various signal-to-noise ratios, ideally for testing ASR system’s robustness to noise.
+For the evaluated models, we selected Whisper-large-v2 as the ASR model for both greedy and beam search methods.
+In our proposed GFD approach, we utilized Mistral Jiang et al.
+(2023) as the language model for English datasets, referring to this configuration as GFD-ASR-EN.
+For Chinese and Cantonese datasets, we integrated Breeze Hsu et al.
+(2024) and designated this setup as GFD-ASR-ZH.
+In addition, we benchmark GER, based on task-activating prompting method Chen et al.
+(2024a); Yang et al.
+(2023b), with Instruction-tuned models including Mistral-Instruct.
+We include two oracle word error rates following previous work Hu et al.
+(2024), where the N-Best Oracle $o_{nb}$ denotes the error rate calculated with the best candidate in the N-Best list, and the Compositional Oracle $o_{cp}$ is the best achievable word error rate using all tokens N-Best the list.
+As to OCR tasks, we utilize TrOCR Li et al.
+(2023a) with Mistral and denote the fused model GFD-OCR-EN.
+We benchmark the models on a wide variety of datasets, including Librispeech Han et al.
+(2019), Medical Figure Eight Inc.
+(2019), ATCO2 Szöke et al.
+(2021), Fleurs Conneau et al.
+(2023), NTUML2021Yang et al.
+(2023c), and FormosaSpeech for ASR; NAF Davis et al.
+(2019) for OCR.
+Librispeech is a collection of corpus from audiobooks with subsets "Clean" and "Other".
+Librispeech-Noise is a noised variant of the original LibriSpeech dataset with different signal-to-noise ratios, which is ideal for testing ASR systems’ robustness to noise.
 Medical dataset contains 8.5 hours of medical conversations with associated symptom tags to each audio-text pairs.
@@
 NAF consists of images from U.S national archives with labelled bounding boxes and annotations, ideal for evaluating OCR performance.
-For all ASR experiments, we report word error rate (WER) for evaluations on English datasets, and mix error rate (MER) for that on Chinese datasets.
+For all ASR experiments, we report Word Error Rate (WER) for evaluations on English datasets and Mixed Error Rate (MER) for those on Chinese datasets.
 For OCR experiments, we report Character Error Rate (CER) and Exact Match (EM).
-For exact match, we require that the recognized texts fully matches the ground truth.
-All GFD fused models can be run on a single A6000 GPU.
-We fix $r=0.2$ in Equation 8 for all our experiments; number of beams equaling 5 and 10 for ASR and OCR experiments, respectively.
-4.2 Conventional speech recognition
-We first verify the efficacy of GFD in conventional speech recognition setting and report results in Table 1.
-We notice that beam search improves consistently upon greedy search, and use beam search as the baseline.
-We found that GFD moderately improves on the Medical and Fleurs-HK dataset.
-In the more challenging NTUML2021, we obtained a state-of-the-art 8.83 mixed error rate, surpassing prior work[49] and even the oracle N-Best score.
-This improvement is mostly preserved even when we narrow the target space to only consider English terms in the dataset (Table 2), making GFD-ASR an elegant solution for code-switched ASR.
-Examples in the Appendix show that improvement can be attributed to the LLM correcting grammatical mistakes and domain terminologies.
-Performance across Librispeech reveals that GFD improves most on moderate noise conditions but diminishes when the noise is too large $(S/R=-5)$.
-In addition, we do not find improvements in the GER setting using non-GER-specific instruct models, consistent with previous work[5].
-As we already patched malformatted outputs with the best beam result, this increase of error rate is mostly attributed to LLM hallucinations, including incorrect deletions.
-We therefore posit that GFD is more robust for incorporating off-the-shelf LLMs on the ASR task than GER.
-Table 1: Results on conventional speech recognition.
-Whisper GER GFD-ASR-EN Oracle
-EN dataset Greedy 5-beams (Baseline) $\text{O}_{\text{nb}}$ $\text{O}_{\text{cp}}$
-Librispeech-Clean 2.30 2.28 3.21 2.29 1.91 1.63
-Librispeech-Other 5.23 4.97 6.57 4.99 4.20 3.43
-Medical 7.30 7.22 7.49 6.74 6.17 5.05
-Librispeech-Noise ($S/R=10$) 3.50 3.12 4.05 3.02 2.32 1.94
-Librispeech-Noise ($S/R=5$) 5.67 5.25 6.00 4.96 4.10 3.28
-Librispeech-Noise ($S/R=0$) 15.23 13.54 14.44 13.27 11.66 9.04
-Librispeech-Noise ($S/R=-5$) 49.09 47.05 47.35 46.98 43.62 33.19
-Whisper GER GFD-ASR-ZH Oracle
-ZH or HK dataset Greedy 5-beams (Baseline) $\text{O}_{\text{nb}}$ $\text{O}_{\text{cp}}$
-Fleurs-HK (Cantonese) 7.49 6.88 7.33 6.23 5.58 4.58
-NTUML2021 11.11 9.97 9.87 8.83 8.88 4.54
-NTUML2021 FormosaSpeech
-Full English
-Whisper (short) 9.97 15.21 -
-SICL (short) [49] 9.5 - -
-GFD-ASR-ZH (short) 8.83 13.51 -
-Whisper (long) 8.40 11.56 27.41
-GFD-ASR-ZH (long) 8.18 11.43 26.72
-Table 2: Comparison results on long-form and short-form.
-4.3 Long-form speech recognition
-As LLMs are capable in attending to long context of sequence and possess multilingual capabilities, we posit that GFD-ASR-ZH can improve on long-form audio transcription.
-We therefore evaluate the long-form transcription performance on NTUML2021 and FormosaSpeech.
-In constrast to the short-form evaluations in Table 1 in NTUML2021, we prepend all previous transcription from contiguous clips in the same lecture as prompts for both ASR and LLM to realize long form transcriptions, but truncate it when necessary for whisper due to context length limitations.
-In the setting of evaluations on contiguous short clips, GFD-ASR-ZH (long) consistently outperforms Whisper (long) in both NTUML2021 and FormosaSpeech, demonstrating that the long-context capability of LLM can be effectively utilized through GFD.
-4.4 Context-aware speech recognition
-Table 3: Results on context-aware ASR task. We set num beams=5 across all our experiments.
 Method Prompt ATCO2 Librispeech FormosaSpeech Medical
 ASR LLM Norm Raw Clean Other
-Whisper No - 47.70 66.44 2.28 4.97 27.42 7.22
+RobustGER, UADF Chen et al. (2024b) No Yes >100 >100 >100 >100 >100 >100
+Clairaudience Liao et al. (2023) Yes - 28.77 - - - - 6.54
+RobustGER Hu et al. (2024) Yes No 34.77 50.58 - - - -
+Whisper No - 47.70 66.44 2.28 4.97 22.33 7.22
 Whisper Yes - 31.34 42.37 - - - 6.24
-Clairaudience[35] Yes - 28.77 - - - - 6.54
-GFD-ASR-EN No Yes 38.75 52.24 2.20 4.61 26.29 6.62
+GFD-ASR-EN No Yes 38.75 52.24 2.20 4.61 20.59 6.62
 GFD-ASR-EN Yes Yes 25.79 32.46 - - - 6.26
-In context-aware speech recognition, we aim to investigate the capability of the fusion model in leveraging contextual information, which is highly desired in real-life contexts, where the incoming speech is expected to be bounded by domain, contain certain specific words, or follow a certain transcribing format.
-We utilize prompting with ASR model and the LLM model (in the GFD setting) to condition on these external hints, in three following settings.
-Domain and Rare Word Prompting: We tested domain-conditioned ASR on the Medical dataset, where the symptom tags are provided along with the speech content.
+Table 3: Results on instruction-aware ASR task.
+All experiments are done with a beam size of 5.
+These experiments is conditioned on a given prompt containing either domain tags (on Medical), rare words (on Librispeech and FormosaSpeech), and complex transcription guidelines (on ATCO2).
+Error rates of over 100 is reported with GER-based methods (Hu et al., 2024; Chen et al., 2024b), as they catastrophically fail on prompt conditioning due to their inability to process instruction prompts beyond the GER prompt.
+4.2 Short-form speech recognition
+We verify the efficacy of GFD in speech recognition setting and report results in Table 1.
+First, we notice that for decoding results without incoporating an LLM, beam search improves consistently upon greedy search.
+Using beam search as the baseline, we observe that GFD moderately improves on the Medical and Fleurs-HK dataset.
+In the more challenging NTUML2021, we obtained a 8.83 mixed error rate, surpassing even the oracle N-Best score.
+Upon examining the benchmarked samples, we attribute the observed improvements to the LLM’s ability to correct English grammatical mistakes and domain-specific terminology within the code-switching context.
+As such, GFD serves as an elegant solution that facilitates the success of code-switched ASR systems.
+Performance across Librispeech demonstrates that GFD offers the most significant enhancement under moderate noise conditions but diminishes when the noise level is too high $(S/R=-5)$.
+In contrast, we do not find improvements in the GER setting using general instruct models, consistent with previous work Chen et al.
+(2024a).
+The increased error rate is primarily attributed to LLM hallucinations, including incorrect deletions.
+Therefore, we posit that GFD is more robust than GER for incorporating off-the-shelf LLMs in the ASR task.
+As demonstrated in Table 2, GFD achieves similar improvements when compared to a specialized GER model, RobustGER Hu et al.
+(2024).
+Analyzing the outputs reveals that the corrected errors are orthogonal: while GER is specifically instructed to adhere to the words in the N-best list, GFD can select words from an exponential search space with intermediate interrogation.
+The combination of RobustGER and GFD yields the best results.
+4.3 Long-form speech recognition
+Whisper (5-beams) GFD-ASR-ZH
+NTUML2021 (long-form) 8.40 8.18
+FormosaSpeech 22.33 20.59
+Table 4: Performance for long-form speech recognition. For NTUML2021, we concatenate all contiguous clips to reconstruct the original lecture for long-form evaluation.
+LLM’s capability to attend to long sequences, makes it an appealing candidate on long-form audio speech recognition.
+Therefore, we evaluate long-form transcription performance on NTUML2021 and FormosaSpeech.
+For NTUML2021, we curate the long dataset by concatenating all contiguous clips to reconstruct the original lecture for long-form evaluation.
+To properly contrast the short-range modeling in Table 1, we prepend all historical transcriptions as prompts for ASR and LLM to realize long-form transcriptions, truncating them when necessary for Whisper due to context length limitations.
+In this setting, GFD-ASR-ZH consistently outperforms Whisper in both NTUML2021 and FormosaSpeech, demonstrating that the long-context capability of LLM can be effectively utilized through GFD (Table 4).
+4.4 Instruction-aware speech recognition
+In instruction-aware speech recognition, we explore GFD’s ability to leverage contextual information, which is crucial in real-life scenarios where speech may be domain-specific, contain rare or critical terms, or require adherence to complex transcription guidelines.
+We employ prompting with ASR and LLM models along with GFD to incorporate these external cues across the respective three settings.
+Domain tag and rare word prompting.
+We tested domain-conditioned ASR on the Medical dataset, where the symptom tags are provided along with the speech content.
 Results show that our GFD method with LLM prompting improves on GFD without prompting, showing prompt sensitiveness of the LLM in the GFD system.
 However, we did not find further improvement upon whisper prompting, compared to double-prompted GFD.
-For verifying rare word prompting capabilities, we used the augmented Librispeech and FormosaSpeech dataset, where a target rare word is mixed with 100 other distractors for each data point [28].
+For verifying rare word prompting capabilities, we used the augmented Librispeech and FormosaSpeech dataset, where a target rare word is mixed with 100 other distractors for each data point Le et al.
+(2021).
 For the FormosaSpeech Dataset, we created the rare words with ChatGPT, and generate distractors with a similar approach using the training set.
 At this scale of distractors, it is unrealistic to prompt on Whisper, as the remaining context length is often insufficient for ASR decoding.
-By prompting on the LLM using GFD, we demonstrated up to 6% WERR over non-prompting methods on the Librispeech dataset, and 1.6% WERR on the FormosaSpeech Dataset.
-Instruction Aware Prompting: We evaluated formatted speech recognition on the ATCO2 dataset, a dataset on air traffic control communications, which has strict regulations on call signs and transcribe formats.
-By incorporating a LLM, we are able to prompt it with over 4000 words of guidelines from an enitre instruction manual [^3] https://www.faa.gov/air_traffic/publications/atpubs/aim_html/chap4_section_2.html.
-a possibility not present with Whisper.
-For Whisper prompting, we include all special call signs and three example sentences extracted from the manual, where we leave details in the Appendix.
-Results in Table 3shows that GFD with ASR and LLM prompting obtains best results with a WERR of 17.7% compared with Whisper in the normalized (ATCO2-Norm) setting, even outperforming Clairaudience[35], a fine-tuned prompt conditioning model.
-We also reported scores without word conversion normalization (ATCO2-Raw), from observations that standard normalization can excessively correct errors that conflicts with the transcribing guidelines.
-In this setting, our gains are even larger, showing instruction following abilites of the GFD system.
+By prompting on the LLM using GFD, we demonstrated up to 7% WERR over non-prompting methods on the Librispeech dataset, and 1.6% WERR on the FormosaSpeech Dataset (Table 3).
+Prompting with Instructions.
+We evaluated formatted speech recognition on the ATCO2 dataset, a dataset on air traffic control communications, which has strict regulations on call signs and transcribe formats.
+By incorporating an LLM, we are able to prompt it with over 4000 words of guidelines from an entire instruction manual [^2] https://www.faa.gov/air_traffic/publications/atpubs/aim_html/chap4_section_2.html, a possibility not present with Whisper.
+For Whisper prompting, we include all special call signs and three example sentences extracted from the manual (Appendix A.2).
+Results in Table 3 show that GFD with ASR and LLM prompting obtains best results with a WERR of 17.7% compared with Whisper in the normalized (ATCO2-Norm) setting, even outperforming Clairaudience Liao et al.
+(2023), a fine-tuned prompt conditioning model.
+GER-based methods Hu et al.
+(2024); Chen et al.
+(2024b) also fall short in this category, due to their inability to process instruction prompts beyond the GER prompt.
+Transcription results completely diverge from the spoken content, causing meaningless error rates exceeding 100%.
+We also reported scores without word conversion normalization (ATCO2-Raw), from observations that standard normalization, such as converting arabic to written numerals, can excessively correct errors that conflicts with the transcribing guidelines.
+In this setting, our improvements are even more pronounced, further demonstrating the instruction-following capabilities of the GFD system.
 4.5 Optical Character Recognition
-Table 4: Evaluation on OCR task
+We use the OCR task as an example to demonstrate that GFD is applicable to auto-regressive scenarios beyond ASR.
+In Table 5, we show that fusing the Mistral LLM to the TrOCR model significantly improves the OCR results on the National Archive Forms dataset by 16.7 % in character error rate reduction and 38.07 % in exact match improvement.
 CER $\downarrow$ Exact Match $\uparrow$
@@
 GFD-OCR 10.55 33.33
-We demonstrate that GFD can be applied to fusing end-to-end OCR model with LLMs, using TrOCR and Mistral [32, 26].
-We experiment the fusion model on a dataset extracted from National Archive Forms dataset [13], denoted NAF-Long.
-The NAF-Long dataset consists of images with text sequences having character lengths between 100 and 200, equivalent to around 25 to 50 text tokens.
-In Table 4, we show that fusing LLM to the OCR model significantly improves the OCR results by 16.7 % in character error rate reduction and 38.07 % in exact match improvement.
-5 Discussion
-Further comparisons with GER.
+Table 5: Evaluation on NAF-Long (an OCR task).
+5 Analysis
+5.1 Further comparisons with GER
 The GFD algorithm relates to GER in that both algorithms perform the selection of output sequences with beam decoding.
 However, they differ in compute execution and the diversity of sample sequences.
-First, in GFD, the LLM works in parallel with the ASR decoder, and thus the computation of per step inference can be executed asynchronously with a load bounded by $\mathcal{O}(Z)+\mathcal{O}(k\cdot\max(S_{ASR},{S_{LLM}}))$, where $Z$ denotes the size of speech encoding, $k$ is the beam size, and $S_{ASR}$ and $S_{LLM}$ are ASR and LLM decoding lengths, respectively.
+First, in GFD, the LLM works in parallel with the ASR decoder, and thus the computation of per step inference can be executed asynchronously with a load bounded by $\mathcal{O}(Z)+\mathcal{O}(k\cdot\max(S_{ASR},{S_{LLM}}))$, where $Z$ denotes the size of speech encoding, $k$ is the beam size, and $S_{ASR}$ and $S_{LLM}$ are ASR and LLM decoding costs of a single token, respectively.
 The LLM decoding complexity expression matches that of Section 3.2, treating the vocabulary size as a constant in this dicussion.
@@
 This means the execution time of GER is bounded by $\mathcal{O}(Z)+\mathcal{O}((k+1)\cdot S_{LLM})+\mathcal{O}(k\cdot S_{ASR})$, where the additional $O(S_{LLM})$ comes from LLM decoding.
-Secondly, in GFD, the searched token space attended by the LLM is at least n_beams * sequence length, whereas in GER, the top-k cutoff of the ASR step does not promote diversity between the candidates, which limits the LLM search space.
+Secondly, in GFD, the searched token space attended by the LLM is at least n_beams times sequence length, whereas in GER, the top-k cutoff of the ASR step does not promote diversity between the candidates, which limits the LLM search space.
 Aside from these differences, GFD and GER are methodologically orthogonal, allowing for their combination in the pursuit of further improvement.
-Limitations. Despite promsing results, we identified three cases where GFD is most susceptible to failure, which is an exciting future direction. We categorize them into ASR errors and LLM errors.
+5.2 Further comparisons with Other Late Fusion Approaches
+Concurrently with our work, step-by-step synchronous late fusion methods that focus on resolving tokenization mismatch have been explored Mittal et al.
+(2024); Hori et al.
+(2025).
+However, these approaches impose constraints on the scoring process, limiting it to specific decoding configurations.
+For example, in SALSA Mittal et al.
+(2024), the LLM sequence is only rescored if it is ending in a UTF-8 character, a criterion designed for non-ascii languages such as Mandarin and Hindi.
+Similarly, Delayed Fusion Hori et al.
+(2025) discards trailing partial tokens from the ASR that do not form complete words during LLM rescoring, which is not easily extensible to languages without explicit word boundaries, such as Mandarin.
+In contrast, our method approaches the problem from the byte sequence level, enabling a more general and flexible rescoring process that supports arbitrary input sequences.
+From an information-theoretic perspective, our approach preserves the maximum available information at each decoding step, avoiding sequence cutoffs or step skipping.
+Empirical results in Table 6 support our hypothesis, with GFD outperforming other settings across English and Mandarin.
+As further detailed in Appendix A.1, the computational overhead introduced by branching through GFD is negligible compared to the cost of a single forward pass of the large language model.
+Method Libri.-Noisy-5 (en) Formosa-Sp. (zh)
+GFD (Ours) 4.96 20.59
+No Branching 4.98 21.29
++ Word Cutoff 5.18 22.33x
++ Char. Cutoff 4.98y 21.50
+Table 6: Word error rates of GFD compared with alternative rescoring methods across English and Mandarin.
+After removing branching, additional constraints can be applied—such as discarding trailing partial tokens, similar to Hori et al.
+(2025), or removing trailing bytes that do not form complete UTF-8 characters, similar to Mittal et al.
+(2024).xThis setting reduces to no rescoring due to the absence of explicit word boundaries.
+yThis setting is identical to no branching in ASCII languages.
+5.3 Error Analysis
+Despite promising results, we identified three cases in which GFD is most susceptible to failure, representing exciting directions for future work.
+We categorize these failure cases into ASR errors and LLM errors.
 ASR proposed errors: When ASR proposes a candidate with high probability that deviates from the phonetic constraint, the LLM may falsely pickup the sequence, inducing an error.
@@
 Semantically activated tokens are tokens that are encoded in similar output embedding space due to semantic similarity.
-We found that these errors are much more common in Chinese and identified some pairs from a top-k analysis of a decoding sequence, e.g., {CJK}UTF8bsmi’剛’ and’才’, ’做(什麼)’ and ’幹(什麼)’ .
+We found that these errors are much more common in Mandarin.
 Time delay activated tokens are proposed tokens that are targets at a later step, where selecting them effectively skips some intermediate tokens.
@@
 6 Conclusion
-The Generative Fusion Decoding (GFD) framework represents a leap forward in fusing auto-regressive text models from heterogeneous token spaces.
-Our experiments confirm GFD’s effectiveness, particularly in ASR tasks.
-For tonal languages like Chinese, it achieves a mixed error rate of 8.83% on the NTUML2021 dataset, surpassing previous benchmarks.
-In context-aware recognition, GFD demonstrated its prowess by a 17.7% Word Error Rate Reduction (WERR) on the ATCO2 dataset when utilizing prompts to both ASR and LLM .
-For OCR tasks, GFD’s fusion with LLMs resulted in a notable 38.06% improvement in exact matches on the NAF-Long dataset, underscoring the framework’s ability to enhance recognition of long text sequences.
-These experimental outcomes highlight GFD’s robustness and adaptability, showcasing its generality to leverage the extensive training and contextual understanding of LLMs to improve text recognition across various domains and challenges.
-We hope that our framework will motivate further research into more fusion techniques that better leverages the existing capabilties of the pre-trained models.
+In summary, we propose Generative Fusion Decoding (GFD), a simple yet effective framework for integrating large language models into ASR and OCR systems through byte-level shallow fusion.
+Our theoretical derivations provide a foundation for this integration, while empirical results demonstrate consistent gains across diverse conditions—including noisy audio, long-form inputs, and instruction-following tasks.
+GFD also compares favorably with prior fusion methods, highlighting its potential as a general-purpose solution.
+These results establish a foundation for future exploration of fusion methods that further exploit the strengths of pre-trained language models.
+Limitations
+The effectiveness of GFD is hindered when LLM selects an ASR token candidate that deviate from the correct phonetic content, leading to hallucinations.
+We provide in our analysis general categories of these errors, for practitioners to be aware of such a risk.
+We also advise users to carefully select the LLM, as the LLM itself may have limitations in its understanding or biases present in its training data.
+If the LLM misinterprets context or generates incorrect predictions, these errors can propagate through the GFD framework, affecting the overall performance.
+Appendix A Appendix
+A.1 Recursive Calculation of the GFD formula
+We have shown the efficient calculation of $\mathbb{P}_{m,approx}(\{B_{l}\},\mathcal{Z}^{(m)})$ with equation (7).
+We now show that the incremental calculation of $\mathbb{P}_{m,approx}(\{B_{l}\},\mathcal{Z}^{(m)})$ from $\mathbb{P}_{m,approx}(\{B_{l-1}\},\mathcal{Z}^{(m)})$ is $O(1)$ in terms of the costly decoder forward.
+Since the model forwarding is only dependent on $\hat{T}_{s}$, and not on alternative tokens, we first derive how the main sequence differs between $B_{l}$ and $B_{l-1}$.
+We denote the main sequence of $B_{l}$ as $\hat{T}_{s}$, and let the main sequence of $B_{l-1}$ as $\hat{T\textquoteright}_{s}$.
+In most scenarios, $\hat{T}_{s}$ is one of the two:
+  • $\hat{T^{\prime}}_{s}$ plus one additional byte token.
+  • The additional byte token merges with previous tokens in $\hat{T^{\prime}}_{s}$, a new token appends a truncated $\hat{T^{\prime}}_{s}$
+In either of the cases, there will only be one additional token on the main path.
+Calculating alternative tokens only requires indexed selection through operations like "mask-select", which is inexpensive compared to the model forward operation.
+[^3] The total FLOPs for the byte algorithm remain under $10^{5}$, which is negligible compared to model forwarding $(>10^{10})$ Therefore, with proper kv-caching on results of $B_{l-1}$, we can efficiently calculate $B_{l}$ to realize GFD.
+A.2 Experimental Details on selecting $r$ in Equation 8
+All GFD fused models are run on a single A6000 GPU.
+For the parameter of $r$ in Equation 8, we conduct grid search of among $[0.1,0.2,0.3,0.4]$ on noisy-librispeech, and selected $r=0.2$.
+We keep $r=0.2$ across all our experiments; while setting the number of beams equal to 5 or 10 for ASR and OCR experiments, respectively.
+A.3 Experimental Details on selecting $k$ in Equation 8
+To maintain reasonable time complexity, we aim to limit the number of rescoring samples to match the number of beams, i.e., num_beams.
+Assume num_beams=5.
+During the expansion phase of beam search, when $k=0$, the text recognition modality generates up to $5\times 5=25$ candidates, which is excessive.
+By strategically selecting $k$ such that the resulting sequence consists only of tokens from sequences prior to the expansion phase, the number of candidates will naturally be capped at the beam size.
+Thus, the optimal value of $k$ corresponds to the length of the last token proposed by the text recognition modality, varying across different beam hypotheses.
+Choosing a larger $k$ results in a loss of information, which is suboptimal.
+A.4 Prompting Details
+Here we list the prompting details of benchmarking.
+Librispeech and noisy-librispeech
+ASR Prompt: (None)
+LLM Prompt: {mdframed} The following is a transcription of a spoken sentence:
+Medical
+ASR Prompt: (None)
+LLM Prompt: {mdframed} The following is a transcription of a spoken sentence:
+Fleurs-HK
+ASR Prompt: {mdframed} (In Chinese) The following is a Traditional Chinese Transcription: LLM Prompt: {mdframed} (In Chinese) The following is a Traditional Chinese Transcription:
+ML Lecture
+ASR Prompt: {mdframed} (In Chinese) Traditional Chinese LLM Prompt: {mdframed} (In Chinese) The following is a Traditional Chinese Transcription, there exists code-switching, and some of the vocabulary is in English.
+Formosa
+ASR Prompt: {mdframed} (In Chinese) The following is a Traditional Chinese Transcription LLM Prompt: {mdframed} (In Chinese) The following is a Traditional Chinese Transcription:
+ATCO2
+ASR Prompt: {mdframed} Alfa Bravo Charlie Delta Echo Foxtrot Golf Hotel India Juliett Kilo Lima Mike November Oscar Papa Quebec Romeo Sierra Tango Uniform Victor Whiskey Xray Yankee Zulu One Two Three Four Five Six Seven Eight Nine Zero
+Dayton radio, November One Two Three Four Five on one two two point two, over Springfield V-O-R, over.
+New York Radio, Mooney Three One One Echo.
+Columbia Ground, Cessna Three One Six Zero Foxtrot, south ramp, I-F-R Memphis.
+LLM Prompt: {mdframed} Section 2.
+Radio Communications Phraseology and Techniques
+1. General
+…(4000 words on call signs and regulations)…
+Generative Error Correction
+We follow Task-Activating Prompting method in Chen et al. (2024a) to create the prompt for Generative Error Correction. {mdframed} User: Do you know Automatic Speech Recognition?
+Assistant: Yes, I do! …
+User: Do you know language model restoring…
+Assistant: Language model restoring is …
+User: Can you generate an example with 5-best list?
+Assistant: Sure! …
+User: Please do the same thing on the following n-best list…